OpenStudy (anonymous):

Create an equation with point-slope form for: passes through (1, 2) with slope of -5/3

OpenStudy (slaaibak):

\[y - y_1 = m (x - x_1)\] Just plug everything in that formula

OpenStudy (slaaibak):

where (1,2) = (x1,y1) and m = -5/3

hero (hero):

Slay bauch

OpenStudy (slaaibak):

sup hero!